what is the simflified answer and the step on how to do the problem (x^5/3 *x^-9/2)^-12/17
Answer by dnanos(83)

(x^5/3 *x^-9/2)^-12/17=[x^(5/3-9/2)]^-12/17=[x^(10/6-27/6)]^-12/17=[x^(-17/6)]^-12/17=x^[(-17/6)*(-12/17)]=x^[(17*12)/(6*17)]=x^(12/6)=x^2